# A priori error analysis of the proximal Galerkin method

### v2

@@ 1. Introduction
This paper presents the first general a priori error analysis of the PG method, which has demonstrated competitive efficacy across a range of problems in applied mathematics, including classical obstacle, contact, and elastoplasticity problems [19]. Prior analyses have focused on specific discretization choices for particular problems or examined the convergence properties of the PG subproblems only after linearization [36, 28]. Instead, the present work provides a foundational advancement, dispensing with analysis of the linearized subproblems and developing a general analytical framework for PG methods applied to quadratic optimization problems in Sobolev Hilbert spaces with pointwise inequality constraints.
To illustrate the flexibility of the proposed framework, we focus on two canonical applications — the obstacle and Signorini problems — demonstrating optimal error convergence rates across a range of finite element discretizations. Key contributions of this work include the first general convergence guarantees with respect to both iteration count and mesh size, as well as the first result establishing mesh-independent iteration complexity of the PG method. These theoretical advances provide rigorous explanations for the method’s empirically observed advantages over well-established methods that preceded PG in the literature, such as the penalty and primal-dual active set methods. We refer the interested reader to [19, 40] for comparisons with these and other popular alternative methods.
@@ 1.2. Notation
where $\tilde{w}$ is the extension by zero of $w$ outside of $\Gamma$; i.e., $\tilde{w}=0$ on $\partial\Omega\backslash\overline{\Gamma}$ and $\tilde{w}=w$ on $\Gamma$. We note that this space is normed by $\|w\|_{\widetilde{H}^{1/2}(\Gamma)}:=\|\tilde{w}\|_{H^{1/2}(\partial\Omega)}$ and that ${H}^{-1/2}(\Gamma)=(\widetilde{H}^{1/2}(\Gamma))^{\prime}$. For non-integer $s$, the notation $H^{s}(\Omega)$ denotes the Sobolev–Slobodeckij spaces [23, Chapter 2]. We use the notation $(\cdot,\cdot)_{\omega}$ to denote the $L^{2}(\omega)$-inner product over measurable $\omega\subset\overline{\Omega}$. If $\omega=\Omega$, we drop the subscript and denote the $L^{2}$-inner product over $\Omega$ by $(\cdot,\cdot)$.
@@ 1.3. General setup
for positive constants $\nu$ and $M$, and $F$ is a bounded linear functional on $V$. The convexity of $K$ implies that the model problem 4 is equivalent to the following variational inequality [16, Theorem 6.1-2]: find $u^{*}\in K$, such that
@@
Owing to the coercivity of $a$, 9 admits a unique solution [10, Theorem 5.6]. Moreover, $E(u^{\ast})\leq E(v)$ for all $v\in K$.
@@ 1.4. The proximal Galerkin method
We present the method here and refer to Section 3 for more details on its derivation. Consider two conforming discrete subspaces $V_{h}\subset V$ and $W_{h}\subset W:=L^{\infty}(\Omega_{d};\mathbb{R}^{m})$. The PG method, given in Algorithm 1, consists of iteratively solving for primal solutions $u_{h}^{k}\in V_{h}$ and latent solutions $\psi_{h}^{k}\in W_{h}$. Note that some form of Newton’s method is usually used to solve each subproblem in practice; see, e.g., the implementations in [20].
@@ 1.5. Main results
  • We provide a general framework for the error analysis of Algorithm 1. This framework shows that the existence and optimality of certain enriching and Fortin operators, defined below, are sufficient to derive error estimates and mesh-independence results; see Theorem 3.3, Theorem 3.6, and Corollary 1.
  • We demonstrate applications of this framework to the analysis of obstacle and Signorini problems; see Section 4 and Section 5. In particular, we construct and prove error rates for the Fortin and enriching operators used in the proposed framework of Section 3.
@@
    for all $\ell\geq 1$ and $h>0$, where the constants $C_{\rm{stab}}$ and $C_{\rm{reg}}$ are independent of $h$, $\{\alpha_{k}\}$, and $\ell$. We first state 15 in Theorem 3.6, where we prove it under general assumptions. We then prove 15 again in Corollaries 2 and 3 for the obstacle and Signorini problems, respectively, by verifying the general assumptions for specific choices of $V_{h}$ and $W_{h}$.
@@ 1.6. Example problems
Korn’s inequality guarantees the coercivity of the bilinear form $a$ over $V$; see, e.g., [25, Theorem 42.9]. In this case, $\lambda^{*}=(\mathsf{C}\epsilon(u^{*})n)\cdot n\in{H}^{-1/2}(\Gamma_{\mathrm{T}})=Q^{\prime}$.
@@
In 22, $\gamma>0$ is a fixed parameter and $\operatorname{proj}$ is the orthogonal projection $H_{0}(\operatorname{div},\Omega)\to\{p\in H_{0}(\operatorname{div},\Omega)\mid\operatorname{div}p=0\}$.
@@
Note that $a$ is coercive over the Hilbert space $V$ due to a Friedrichs’ inequality; see, e.g., [18, Lemma 2.8].
@@ 1.7. Closed observation maps: A conjecture
Some important problems with feasible sets of the form 5 do not fit into the general setup described in Section 1.3 because the observation map $B$ does not map onto a dense subset of $L^{2}(\Omega_{d};\mathbb{R}^{m})$. For instance, consider the classical elastoplastic torsion problem [48, 9], which involves minimizing the Dirichlet energy 16 over a feasible set with gradient constraints, such as
@@
Note that this set is recovered from 5 by setting $C$ to be the closed unit ball in $\mathbb{R}^{n}$, $V=H^{1}_{0}(\Omega)$, $B=\nabla$, and $\Omega_{d}=\Omega$. For further details, see [19, Example 6]. If $n\geq 2$, then $\operatorname{im}B$ is a closed proper subspace of $L^{2}(\Omega;\mathbb{R}^{n})$.
@@ 2.1. Legendre functions
In the sequel, we tacitly assume a supercoercivity property of the map $R(x,\cdot)$; namely, $R(x,y)/|y|\rightarrow\infty$ as $|y|\rightarrow\infty$ for a.e. $x\in\Omega$. This ensures that $R^{*}(x,\cdot)$ is well-defined and continuously differentiable over all of $\mathbb{R}^{m}$ [6, Proposition 2.16]; see also [42, Corollary 13.3.1]. Likewise, we can conclude that $\nabla\mathcal{R}^{*}$ is continuous on $L^{\infty}(\Omega_{d};\mathbb{R}^{m})$. We now utilize the following relation, first demonstrated for Legendre functions in [41]:
@@
In turn, we conclude that $\operatorname{dom}(\nabla\mathcal{R})=\operatorname{im}(\nabla\mathcal{R}^{*})\subset\mathcal{O}$, implying $\nabla\mathcal{R}^{*}(\psi)(x)\in\operatorname{int}C(x)$ f.a.e. $x\in\Omega_{d}$. We provide three explicit examples in Examples 2.1, 2.2 and 2.3 below, and we refer to [19, Table 1] for more.
Identity 26 also allows us to define a latent representation of every observable $Bu\in\operatorname{dom}(\nabla\mathcal{R})$; namely,
@@
In the subsequent sections, we will make use of the following identity, which can be derived by directly expressing the convex conjugate $R^{*}(x,z)$ given in 25 by
@@
The continuity of $\mathcal{R}$ on $\mathcal{O}$ (Assumption 1) follows from [36, Theorem 4.1]. A simple computation shows that $R^{*}(x,z)=\exp(z)+\phi(x)z$ with
@@
if $\underline{u}(x)\leq y\leq\overline{u}(x)$ and $R(x,y)=+\infty$ otherwise. The corresponding superposition operator $\mathcal{R}$ is continuous on $\mathcal{O}$; i.e., it satisfies Assumption 1 [36, Lemma 3.2]. Here, $\nabla\mathcal{R}(o)=\ln(o-\underline{u})-\ln(\overline{u}-o)$ with
@@
With 26, we derive that
@@
For Euclidean norm constrains $|Bu|\leq\gamma$, cf. 24, one can select the Hellinger entropy
@@
This entropy has been used in the PG framework for gradient norm constraints, see Section 1.7 and [19, Example 6].
@@ 2.2. Bregman divergences
The proof can be found in [2], see also [6, Theorem 3.9]. We also recall the three points identity [13, Lemma 3.1]:
@@ 2.3. Deriving the proximal Galerkin method
The algorithm 37 leverages the Legendre function $\mathcal{R}$ to adaptively regularize 4, resulting in iterates $u^{k}$ that converge at a controllable speed to the global minimizer $u^{\ast}$. The message behind the LVPP reformulation is that the subproblems in 37 are easy to discretize and solve if the latent variable $\psi$ in 27 is incorporated.
Formally, choosing $R(x,\cdot)$ with singular derivatives at $\partial C(x)$ f.a.e. $x\in\Omega_{d}$ (cf. Section 2.1), one expects that $Bu^{k}\in\operatorname{im}\nabla\mathcal{R}^{*}$; in particular, $Bu^{k}(x)\in\mathrm{int}\,C(x)$ f.a.e. $x\in\Omega_{d}$. Likewise, the solutions $u^{k}$ of the regularized subproblems 37 are characterized by variational equations:
@@
for all $v\in V$. Introducing the latent variables $\psi^{k}=\nabla\mathcal{R}(Bu^{k})$ (i.e., $Bu^{k}=\nabla\mathcal{R}^{*}(\psi^{k})$ by 26) to rewrite the resulting equations in saddle-point form yields the LVPP algorithm, and discretizing the resulting saddle-point problems leads to the PG method [19]. We refer the reader to [36] for a detailed derivation of LVPP for the obstacle problem and to [28, Section 3] for a brief summary.
@@
and $k=1,\ldots$ Note that from 39b, the latent variables $\psi^{k}$ satisfy a crucial identity: $Bu^{k}=\nabla\mathcal{R}^{*}(\psi^{k})\in\mathcal{O}$. Proving that 39 is well-posed is generally a challenging task. To date, it has only been accomplished for the obstacle problem (cf. Example 1.1) under suitable regularity assumptions on the problem data [36]. Fortunately, in the a priori error analysis that follows, we do not rely on the continuous-level subproblems 39 in any way. Instead, we focus solely on the discretized subproblems 10.
@@ 3. Framework for the analysis of the proximal Galerkin method
This section presents our main results in a general framework. We first introduce a compatible subspace condition to demonstrate that the PG method is well-defined. We then show that PG is endowed with a convenient energy decay property, leading to best approximation results and error convergence rates if a so-called enriching operator exists. The assumptions of this section are verified for the obstacle and Signorini problems in Section 4 and Section 5, respectively.

### v3

@@ 1. Introduction
This paper presents general a priori error analysis of the PG method, which has demonstrated competitive efficacy across a range of problems in applied mathematics, including classical obstacle, contact, and elastoplasticity problems [21]. Prior analyses have focused on specific discretization choices for particular problems or examined the convergence properties of the PG subproblems only after linearization [40, 30]. Instead, the present work provides a foundational advancement, dispensing with the analysis of linearized subproblems to arrive at a general analytical framework for PG methods applied to quadratic optimization problems in Sobolev Hilbert spaces with pointwise inequality constraints.
To illustrate the flexibility of the proposed framework, we focus on two canonical applications: the obstacle and Signorini problems. Established approaches for these problems include the quadratic penalty method [5] and “first-discretize-then-optimize” strategies such as the primal-dual active set [36] and augmented Lagrangian methods [33]. Each of these techniques faces its own distinct challenges: penalty methods require scaling the penalty parameter inversely with the mesh size to maintain accuracy [48], while active set and augmented Lagrangian solvers can exhibit mesh-dependent behavior, which can be mitigated via multigrid techniques [34, 12]. There is a vast literature on the error analysis of the aforementioned approaches with various spatial discretization choices. However, we note that the optimization error is often excluded from the analysis when “first-discretize-then-optimize” approaches are employed. For an overview, we refer the interested reader to the textbooks [6, Chapter 5] and [16, Chapters 5 and 6] for the obstacle and Signorini problems, respectively.
In this work, we provide general convergence guarantees for the PG framework with respect to both the number of iterations and the mesh size, as well as the first result establishing its mesh-independent iteration complexity. In particular, we emphasize that PG requires no parameters to be sent to singular limits and is asymptotically mesh-independent without multigrid. We refer the interested reader to [21, 44] for numerical experiments comparing PG to primal-dual active set and other popular alternatives.
@@ 1.2. Notation
where $\tilde{w}$ is the extension by zero of $w$ outside of $\Gamma$; i.e., $\tilde{w}=0$ on $\partial\Omega\backslash\overline{\Gamma}$ and $\tilde{w}=w$ on $\Gamma$. We note that this space is normed by $\|w\|_{\widetilde{H}^{1/2}(\Gamma)}:=\|\tilde{w}\|_{H^{1/2}(\partial\Omega)}$ and that ${H}^{-1/2}(\Gamma)=(\widetilde{H}^{1/2}(\Gamma))^{\prime}$. For non-integer $s$, the notation $H^{s}(\Omega)$ denotes the Sobolev–Slobodeckij spaces [25, Chapter 2]. We use the notation $(\cdot,\cdot)_{\omega}$ to denote the $L^{2}(\omega)$-inner product over measurable $\omega\subset\overline{\Omega}$. If $\omega=\Omega$, we drop the subscript and denote the $L^{2}$-inner product over $\Omega$ by $(\cdot,\cdot)$.
@@ 1.3. General setup
for positive constants $\nu$ and $M$, and $F$ is a bounded linear functional on $V$. The convexity of $K$ implies that the model problem 4 is equivalent to the following variational inequality [18, Theorem 6.1-2]: find $u^{*}\in K$, such that
@@
Owing to the coercivity of $a$, 9 admits a unique solution [11, Theorem 5.6]. Moreover, $E(u^{\ast})\leq E(v)$ for all $v\in K$.
@@ 1.4. The proximal Galerkin method
We present the method here and refer to Section 3 for more details on its derivation. Consider two conforming discrete subspaces $V_{h}\subset V$ and $W_{h}\subset W:=L^{\infty}(\Omega_{d};\mathbb{R}^{m})$. The PG method, given in Algorithm 1, consists of iteratively solving for primal solutions $u_{h}^{k}\in V_{h}$ and latent solutions $\psi_{h}^{k}\in W_{h}$. For the derivation of Algorithm 1 and precise definition of $\mathcal{R}^{*}$, we refer to Section 2.1 and Section 2.3, respectively. Note that some form of Newton’s method is usually used to solve each subproblem in practice; see, e.g., the implementations in [22].
@@
We end this Section with two remarks on the computational aspects of Algorithm 1.
Remark 1 (Equivalent formulation).
We note that (10) can be equivalently reformulated with the variables $(u_{h}^{k},\lambda_{h}^{k})\in V_{h}\times W_{h}$ satisfying:
(15a) $a(u^{k}_{h},v_{h})-b(v_{h},\lambda_{h}^{k})$ $=F(v_{h})$  $~\text{ for all }v_{h}\in V_{h},$
(15b) $b(u^{k}_{h},w_{h})-\left(\nabla\mathcal{R}^{*}(-\alpha^{k}\lambda_{h}^{k}+\psi_{h}^{k-1}),w_{h}\right)_{\Omega_{d}}$ $=0$  $~\text{ for all }w_{h}\in W_{h},$
where $\psi_{h}^{k}$ is updated via the formula
(15c) $\psi_{h}^{k}=-\alpha^{k}\lambda_{h}^{k}+\psi_{h}^{k-1}.$
The formulation (15) is more favorable from a computational perspective since, as we show in Theorems 3.3 and A.1, the variable $\lambda_{h}^{k}$ converges to $\lambda^{*}$ and remains bounded with $k$ and $h$ in the $Q^{\prime}$ norm. This is in contrast to the variable $\psi_{h}^{k}$; see Theorem A.1. Having mentioned this, we note that the $(u_{h}^{k},\psi_{h}^{k})$ formulation of (10) has been successfully implemented in many previous works; see the extensive selection of numerical experiments in [40, 21, 44, 30]. Recognizing that the two formulations are mathematically equivalent, we present the analysis of (10) because it appears to yield simpler arguments.
Remark 2 (Solving the nonlinear subproblems).
The nonlinear subproblems, 10 or, equivalently, 15, can be solved with Newton’s method initialized with the solution of the previous proximal subproblem. In this case, we note that an aggressive update strategy, $\alpha^{k}\gg\alpha^{k-1}$, $k=1,2,\ldots$, may lead to numerical instabilities and, therefore, a large number of linear solves (i.e., Newton steps). In turn, efficiently solving the nonlinear subproblems requires balancing the order of convergence 16 of the proximal (outer) loop with the local stability of Newton’s method (inner loop). In previous work [40, 21], taking $\alpha^{k}=\min(c\alpha^{k-1},C)$, with $c\geq 1,C\gg 1$, has been found to be an effective choice leading to superlinear convergence in most model problems. To increase stability, one can always adaptively decrease (i.e., backtrack) $\alpha^{k}$ whenever the Newton solver struggles. Developing optimal reconditioners for the linear systems arising from Newton’s method is part of ongoing work and is not the focus of this paper. However, we note that block-diagonal preconditioners have been applied efficiently in [44]. Furthermore, reformulating 10 as a first-order system of equations and hybridizing it via facet variables yields symmetric positive-definite stiffness matrices at each Newton step [30], enabling the use of many well-known and efficient preconditioning strategies.
@@ 1.5. Main results
  • We provide a general framework for the error analysis of Algorithm 1. This framework shows that the existence and optimality of certain reconstruction and Fortin operators, defined below, are sufficient to derive error estimates and mesh-independence results; see Theorem 3.3, Theorem 3.6, and Corollary 1.
  • We demonstrate applications of this framework to the analysis of obstacle and Signorini problems; see Section 4 and Section 5. In particular, we construct and prove error rates for the Fortin and reconstruction operators used in the proposed framework of Section 3.
@@
    for all $\ell\geq 1$ and $h>0$, where the constants $C_{\rm{stab}}$ and $C_{\rm{reg}}$ are independent of $h$, $\{\alpha_{k}\}$, and $\ell$. We first state 16 in Theorem 3.6, where we prove it under general assumptions. We then prove 16 again in Corollaries 2 and 3 for the obstacle and Signorini problems, respectively, by verifying the general assumptions for specific choices of $V_{h}$ and $W_{h}$.
@@ 1.6. Example problems
Korn’s inequality guarantees the coercivity of the bilinear form $a$ over $V$; see, e.g., [27, Theorem 42.9]. In this case, $\lambda^{*}=(\mathsf{C}\epsilon(u^{*})n)\cdot n\in{H}^{-1/2}(\Gamma_{\mathrm{T}})=Q^{\prime}$.
@@
In 23, $\gamma>0$ is a fixed parameter and $\operatorname{proj}$ is the orthogonal projection $H_{0}(\operatorname{div},\Omega)\to\{p\in H_{0}(\operatorname{div},\Omega)\mid\operatorname{div}p=0\}$.
@@
Note that $a$ is coercive over the Hilbert space $V$ due to a Friedrichs’ inequality; see, e.g., [20, Lemma 2.8].
@@ 1.7. Closed observation maps: A conjecture
Some important problems with feasible sets of the form 5 do not fit into the general setup described in Section 1.3 because the observation map $B$ does not map onto a dense subset of $L^{2}(\Omega_{d};\mathbb{R}^{m})$. For instance, consider the classical elastoplastic torsion problem [52, 10], which involves minimizing the Dirichlet energy 17 over a feasible set with gradient constraints, such as
@@
Note that this set is recovered from 5 by setting $C$ to be the closed unit ball in $\mathbb{R}^{n}$, $V=H^{1}_{0}(\Omega)$, $B=\nabla$, and $\Omega_{d}=\Omega$. For further details, see [21, Example 6]. If $n\geq 2$, then $\operatorname{im}B$ is a closed proper subspace of $L^{2}(\Omega;\mathbb{R}^{n})$.
@@ 2.1. Legendre functions
In the sequel, we tacitly assume a supercoercivity property of the map $R(x,\cdot)$; namely, $R(x,y)/|y|\rightarrow\infty$ as $|y|\rightarrow\infty$ for a.e. $x\in\Omega$. This ensures that $R^{*}(x,\cdot)$ is well-defined and continuously differentiable over all of $\mathbb{R}^{m}$ [7, Proposition 2.16]; see also [46, Corollary 13.3.1]. Likewise, we can conclude that $\nabla\mathcal{R}^{*}$ is continuous on $L^{\infty}(\Omega_{d};\mathbb{R}^{m})$. We now utilize the following relation, first demonstrated for Legendre functions in [45]:
@@
In turn, we conclude that $\operatorname{dom}(\nabla\mathcal{R})=\operatorname{im}(\nabla\mathcal{R}^{*})\subset\mathcal{O}$, implying $\nabla\mathcal{R}^{*}(\psi)(x)\in\operatorname{int}C(x)$ f.a.e. $x\in\Omega_{d}$. We provide three explicit examples in Examples 2.1, 2.2 and 2.3 below, and we refer to [21, Table 1] for more.
Identity 27 also allows us to define a latent representation of every observable $Bu\in\operatorname{dom}(\nabla\mathcal{R})$; namely,
@@
In the subsequent sections, we will make use of the following identity, which can be derived by directly expressing the convex conjugate $R^{*}(x,z)$ given in 26 by
@@
The continuity of $\mathcal{R}$ on $\mathcal{O}$ (Assumption 1) follows from [40, Theorem 4.1]. A simple computation shows that $R^{*}(x,z)=\exp(z)+\phi(x)z$ with
@@
if $\underline{u}(x)\leq y\leq\overline{u}(x)$ and $R(x,y)=+\infty$ otherwise. The corresponding superposition operator $\mathcal{R}$ is continuous on $\mathcal{O}$; i.e., it satisfies Assumption 1 [40, Lemma 3.2]. Here, $\nabla\mathcal{R}(o)=\ln(o-\underline{u})-\ln(\overline{u}-o)$ with
@@
With 27, we derive that
@@
For Euclidean norm constrains $|Bu|\leq\gamma$, cf. 25, one can select the Hellinger entropy
@@
This entropy has been used in the PG framework for gradient norm constraints, see Section 1.7 and [21, Example 6].
@@ 2.2. Bregman divergences
The proof can be found in [2], see also [7, Theorem 3.9]. We also recall the three points identity [15, Lemma 3.1]:
@@ 2.3. Deriving the proximal Galerkin method
The algorithm 38 leverages the Legendre function $\mathcal{R}$ to adaptively regularize 4, resulting in iterates $u^{k}$ that converge at a controllable speed to the global minimizer $u^{\ast}$. The message behind the LVPP reformulation is that the subproblems in 38 are easy to discretize and solve if the latent variable $\psi$ in 28 is incorporated.
Formally, choosing $R(x,\cdot)$ with singular derivatives at $\partial C(x)$ f.a.e. $x\in\Omega_{d}$ (cf. Section 2.1), one expects that $Bu^{k}\in\operatorname{im}\nabla\mathcal{R}^{*}$; in particular, $Bu^{k}(x)\in\mathrm{int}\,C(x)$ f.a.e. $x\in\Omega_{d}$. Likewise, the solutions $u^{k}$ of the regularized subproblems 38 are characterized by variational equations:
@@
for all $v\in V$. Introducing the latent variables $\psi^{k}=\nabla\mathcal{R}(Bu^{k})$ (i.e., $Bu^{k}=\nabla\mathcal{R}^{*}(\psi^{k})$ by 27) to rewrite the resulting equations in saddle-point form yields the LVPP algorithm, and discretizing the resulting saddle-point problems leads to the PG method [21]. We refer the reader to [40] for a detailed derivation of LVPP for the obstacle problem and to [30, Section 3] for a brief summary.
@@
and $k=1,\ldots$ Note that from 40b, the latent variables $\psi^{k}$ satisfy a crucial identity: $Bu^{k}=\nabla\mathcal{R}^{*}(\psi^{k})\in\mathcal{O}$. Proving that 40 is well-posed is generally a challenging task. To date, it has only been accomplished for the obstacle problem (cf. Example 1.1) under suitable regularity assumptions on the problem data [40]. The main difficulty in proving a general well-posedness result for (40) is in identifying the correct functional setting that typically involves a non-reflexive Banach space $W$. For instance, for the obstacle problem (cf. Example 1.1), it is known that $W=L^{\infty}(\Omega)$ is the correct choice under suitable regularity assumptions on the problem data [40]. Fortunately, in the a priori error analysis that follows, we do not rely on the continuous-level subproblems 40 in any way. Instead, we focus solely on the discretized subproblems 10.
@@ 3. Framework for the analysis of the proximal Galerkin method
This section presents our main results in a general framework. We first introduce a compatible subspace condition to demonstrate that the PG method is well-defined. We then show that PG is endowed with a convenient energy decay property, leading to best approximation results and error convergence rates if a so-called reconstruction operator exists. The assumptions of this section are verified for the obstacle and Signorini problems in Section 4 and Section 5, respectively.